a. Two point charges arranged horizontally. A charge, qì--25.0 μC, is placed at x-0.220 m and a second charge q2 placed at x-0.340 m. A third point charge, q3-+8.40 uC is placed at the origin. The net electrostatic force on q2 is +27.0 N y. What is the charge on qz?
Answer #1

The net force on q2 will be:

Fnet-2 = k q1q2/r12^2 + k q2q3/r23^2

27 = 9 x 10^9 x q2 x 10^-6(25/0.12^2 + 8.4/0.34^2)

27 = 9 x 10^3 x q2 (1808.78)

q2 = 27/(9 x 10^3 x 1808.78) = 1.66 x 10^-6 C

q2 = 1.66 x 10^-6 C

[check, F = 9 x 10^9 x 1.66 x 10^-6 (25/0.12^2 + 8.4/0.34^2) = 27 N]